# E-values, Multiple Testing and Beyond

### v1

@@ C.2 Additional numerical results for structure adaptive multiple testing
The results for $a_{f}=0.5$ are presented in Figure C.1. When the signal is sparse, and the covariate is less informative, slight FDR inflation is observed in SABHA and the two versions of IHW. eBH_GBC has the highest power, followed by SABHA and the two versions of IHW. AdaPT, on the other hand, shows a power loss when compared to the BH procedure. However, as the covariate becomes more informative, all structure adaptive methods outperform the BH procedure, and eBH_GBC has the most true discoveries when the signal is sparse. Furthermore, when the signal becomes dense, eBH_GBC, AdaPT, and SABHA have similar performance in power.
@@
Figure C.2 shows the results for $a_{f}=0$, i.e., the alternative p-value distribution is independent of the covariates. In this case, AdaPT performs the best, followed by eBG_GBC and SABHA, which dominate IHW and the BH procedure.
Table C.1: Parameter settings for the case of $G=4$. Here, $n_{g}$ represents the number of hypotheses for the $g$th group; $n_{ga}$ denotes the number of non-null hypotheses in the $g$th group. The parameter $u_{g}$ is the upper bound of the uniform distribution for the p-values under the alternatives for the $g$th group.
Setting Target FDR level $n_{1}$ $n_{1a}$ $p_{1}$ $n_{2}$ $n_{2a}$ $p_{2}$ $n_{3}$ $n_{3a}$ $p_{3}$ $n_{4}$ $n_{4a}$ $p_{4}$
Setting F1 0.05 100 20 $1.5\times 10^{-3}$ 100 20 $1.5\times 10^{-3}$ 1000 20 $1.5\times 10^{-3}$ 1000 20 $1.5\times 10^{-3}$
Setting F2 0.05 100 1 $5\times 10^{-4}$ 100 20 $2\times 10^{-3}$ 100 20 $3\times 10^{-3}$ 100 20 $4\times 10^{-3}$
Setting F3 0.2 50 2 $1\times 10^{-3}$ 100 2 $1\times 10^{-3}$ 50 4 $2\times 10^{-3}$ 100 4 $2\times 10^{-3}$
[…]
BC_Com 0.698 0.697 0.698 0.698 0.698 0.042 0.006 0.006 0.072 0.073
BC_Sep 0.579 0.906 0.893 0.245 0.273 0.064 0.048 0.045 0.025 0.03
eBH_1 0.022 0 0 0.045 0.045 0.006 0 0 0.005 0.006
eBH_2 0.051 0.051 0.051 0.051 0.05 0.004 0.003 0.002 0.005 0.006
eBH_Ada 0.22 0.313 0.313 0.118 0.137 0.021 0.016 0.015 0.012 0.015
Table C.1: Parameter settings for the case of $G=4$. Here, $n_{g}$ represents the number of hypotheses for the $g$th group; $n_{ga}$ denotes the number of non-null hypotheses in the $g$th group. The parameter $u_{g}$ is the upper bound of the uniform distribution for the p-values under the alternatives for the $g$th group.
Table C.2: FDR and power for Setting F1, where the nominal FDR level is 5%.
Table C.3: FDR and power for Setting F2, where the nominal FDR level is 5%.
[…]
BC_Com 0.93 0.947 0.94 0.932 0.919 0.044 0.312 0.032 0.03 0.031
BC_Sep 0.788 0 0.863 0.794 0.747 0.055 0 0.044 0.043 0.041
eBH_1 0.001 0 0.001 0.001 0.001 0 0 0 0 0
eBH_2 0.001 0 0.001 0.001 0.001 0 0 0 0 0
eBH_Ada 0.541 0 0.56 0.551 0.538 0.032 0 0.029 0.03 0.03
[…]
BC_Com 0.955 0.967 0.968 0.948 0.95 0.174 0.147 0.236 0.095 0.151
BC_Sep 0.393 0.149 0.139 0.528 0.507 0.346 0.1 0.094 0.173 0.171
eBH_1 0.002 0.001 0.003 0.001 0.003 0.002 0.001 0.002 0 0.002
eBH_2 0.007 0.005 0.004 0.008 0.008 0.004 0.003 0.003 0.003 0.003
eBH_Ada 0.038 0.012 0.006 0.053 0.051 0.021 0.008 0.004 0.016 0.018
Table C.3: FDR and power for Setting F2, where the nominal FDR level is 5%.
Table C.4: FDR and power for Setting F3, where the nominal FDR level is 20%.
Figure C.1: Empirical FDR and power for $a_{f}=0.5$.
Figure C.2: Empirical FDR and power for $a_{f}=0$.
wzAYj8cwTZPwPH9/sVg8PXweDAauqqr2cDjEer1GJBLBZDJBs9mE4zjwfZ85lAGg2+06hmGgXq+j3+/DsixYlgVN03a9Xu8jgCNCyIegIAgx13Vfd7vdu+FweG8YRkjXdWy329+dTgeSJD3ieZ7RNO0VAXAPwDEAO5VKndi2fWrb9jWl9Esul6PZbDY9Go1OZ7PZ9z/lyuD3OozU2wAAAABJRU5ErkJggg==" alt="[LOGO]">

### v2

@@ G.2 Additional Numerical Results for Structure-Adaptive Multiple Testing
The results for $a_{f}=0.5$ are presented in Figure G.2. When the signal is sparse and the covariate is less informative, slight FDR inflation is observed in SABHA and the two versions of IHW. eBH_FBC has the highest power, followed by SABHA and the two versions of IHW. AdaPT, on the other hand, shows a power loss when compared to the BH procedure. However, as the covariate becomes more informative, all structure adaptive methods outperform the BH procedure, and eBH_FBC has the most true discoveries when the signal is sparse. Furthermore, when the signal becomes dense, eBH_FBC, AdaPT, and SABHA have similar performance in power.
@@
Figure G.3 shows the results for $a_{f}=0$, i.e., the alternative p-value distribution is independent of the covariates. In this case, AdaPT performs the best, followed by eBG_FBC and SABHA, which dominate IHW and the BH procedure.
Figure G.1: Variance for empirical FDR and power when $a_{f}=1$. Signal sparsity is controlled by setting $a_{0}\in\{3.5,2.5,1.5\}$, giving rise to sparse, moderate, and dense alternatives, respectively. Covariate informativeness is tuned via $a_{1}\in\{1.5,2,2.5\}$, corresponding to weak, moderate, and strong auxiliary signals.
Figure G.2: Empirical FDR and power for $a_{f}=0.5$. Signal sparsity is controlled by setting $a_{0}\in\{3.5,2.5,1.5\}$, giving rise to sparse, moderate, and dense alternatives, respectively. Covariate informativeness is tuned via $a_{1}\in\{1.5,2,2.5\}$, corresponding to weak, moderate, and strong auxiliary signals.
Figure G.3: Empirical FDR and power for $a_{f}=0$. Signal sparsity is controlled by setting $a_{0}\in\{3.5,2.5,1.5\}$, giving rise to sparse, moderate, and dense alternatives, respectively. Covariate informativeness is tuned via $a_{1}\in\{1.5,2,2.5\}$, corresponding to weak, moderate, and strong auxiliary signals.
method BH IHW_Storey IHW_betamix AdaPT SABHA eBH_FBC
time in seconds $3\times 10^{-4}$ $2.3\times 10^{-3}$ $1.9\times 10^{-3}$ 4.24 $6\times 10^{-4}$ 0.45
Table G.1: The average running time for each method in 100 simulations.